Prove that the following two equations represent the simple harmonic motion-

X = A cosωt + B sinωt

X = A eiωt


1
Expert's answer
2021-05-11T09:08:00-0400

The equation of harmonic oscillations is "x(t)=A'\sin(\omega t+\phi)" or "x(t)=A'\cos(\omega t+\phi)" . So,


a) "x(t)=A'\sin(\omega t+\phi)=A'(\sin\omega t\cdot\ \cos\phi+\cos\omega t\cdot \sin\phi)="


"=A'\sin\omega t\cdot\ \cos\phi+A'\cos\omega t\cdot \sin\phi=B\sin\omega t+A\cos\omega t="


"=A\cos\omega t+B\sin\omega t" . Proved



b) For the simple harmonic motion


"\frac{d^2x(t)}{dt^2}+x(t)\omega^2=0"


"x(t)=Ae^{i\omega t}"


"-A\omega^2e^{i\omega t}+Ae^{i\omega t}\omega^2=0" . Proved